OpenStudy (anonymous):

5s^2-4s-6 factor the trinomial

OpenStudy (anonymous):

I gocha

OpenStudy (anonymous):

It is going to become x(5x-4)=6

OpenStudy (anonymous):

You always want to simplify right away and then go from thier

OpenStudy (gorv):

s=(4+(136)^1/2)/10 s=(4-(136)^1/2)/10

OpenStudy (anonymous):

Nooooo, you will divide by 5 and then add 6/5 to each side